Fermi National Laboratory
 

Accelerator Update for May 4, 2007 - May 7, 2007

 

Friday May 4

The day shift began with Operations monitoring Tevatron (TeV) store 5409 (53.34E30) and the Recycler stash (248.1E10), with the Antiproton Source (Pbar) stacking (23.2E10), with the Linac, Booster, and Main Injector (MI) operating normally, with NuMI, MiniBooNE, and the Meson MTest experiments taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) using the colliding beams.

RF experts reported that they repaired a MI RF station (MIRF14) at 11:08 AM.

At 11:30 AM, a Linac Klystron RF station (KRF3) tripped off. It reset.

At 12:04 PM, LRF3 tripped off. It reset.

At 12:29 PM, a Booster RF station (BRF10) tripped off. It reset.

Operations terminated store 5409 at 1:31 PM.

Operations put the TeV into shot setup at 1:47 PM.

At 3:17 PM, KRF4 tripped off. It reset.

Operations reported at 3:58 PM that the Recycler stashed antiprotons once during the day shift.

At 4:15 PM, LRF2 tripped off. It reset.

Operations established store 5410 at 4:24 PM with an initial luminosity of 187.48E30.

At 5:19 PM, KRF4 tripped off. It reset.

At 6:43 PM, the NuMI horn tripped off on a current strip ground fault. It reset.

At 10:22 PM, the #4 helium cold compressor at TeV sector E0 tripped off due to a trip of its cooling water pump. Compressor #1 also tripped off. Operations contacted the cryo coordinator. Over the next hour compressors 1, 2, & 3 also tripped off due to water pump trips. At the end of all this, Operations reported that compressors 2 & 3 were disconnected and had their cooling water valved out. Per instruction, operators opened the cross-connect valve for the E0-EA pond water. E0 compressors 1 & 4 were operating normally. The Duty Electrician reported that the E0 water pump was grounded. This would be repaired later.

At 10:56 PM, BRF2 tripped off. It reset.

Operations reported at midnight that the Recycler stashed three times during the evening shift.

Saturday May 5

The midnight shift began with Operations monitoring store 5410 (79.18E30) and the stash (167.9E10), with Pbar stacking (11.5E10), and with NuMI and MiniBooNE taking beam.

At 12:55 AM, KRF3 tripped off. It reset.

At 5:34 AM, LRF3 tripped off for the first of three times in 15 minutes. It reset.

At 5:47 AM, KRF4 tripped off. It reset.

Operations established beam to MTest at 5:59 AM.

Operations reported at 7:57 AM that the Recycler stashed antiprotons twice during the midnight shift.

At 8:26 AM, LRF3 tripped off. It reset.

At 9:19 AM, Operations reported having a problem with beam closure during the last stash.

At 10:23 AM, a Pbar Debuncher RF station (RF2) tripped off. It reset.

At 10:29 AM, KRF3 tripped off. It reset.

At 12:23 PM, BRF11 tripped off. It reset.

At 12:33 PM, KRF4 tripped off. It reset.

At 1:09 PM, a Recycler power supply (VDPA1) tripped off during a stash and a transfer was lost. Operators found a small LCW leak in the supply and valved it off. Operations called in a power supply expert.

A TeV expert began a Tevatron Electron Lens (TEL1) study at 1:50 PM.

A power supply expert reported at 4:49 PM, that VDPA1 had two soaked transformers.

At 5:33 PM. LRF3 tripped off. It reset.

The MTest experimenters reported at 6:16 PM that they were through with the beam for the night.

The power supply expert working on the Recycler VDPA1 supply reported at 6:47 PM that he replaced a SCR heatsink LCW hose, a firing board, and the ground fault detector/voltage divider panel. He disconnected the delta transformer and the overcurrent module, which were both shorted out. The supply was tested and up and running by 6:48 PM.

Operations terminated store 5410 at 7:04 PM.

Operations put the TeV into shot setup at 7:48 PM.

At 8:02 PM, LRF3 tripped off. It reset.

Operations reported at 8:06 PM that the Recycler stashed twice during the last 12 hours. Accelerator status: the Pbar stack was at 60.5E10 and the stash was at 285.6E10.

At 8:31 PM, the 750 KeV area Buncher tripped off. It reset.

At 8:44 PM, KRF4 tripped off. It reset.

Operations established store 5411 at 9:42 PM with an initial luminosity of 188.87E30.

At 11:13 PM, LRF3 tripped off for the first of two times in 11 minutes. It reset.

Sunday May 6

The midnight shift began with Operations monitoring store 5411 the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

At 12:07 AM, LRF3 tripped off for the first of four times in 90 minutes. It reset.

At 2 AM, LRF3 tripped off for the first of four times over the next five and half hours. It reset.

MTest resumed taking beam at 5:59 AM.

Operations reported at 7:45 AM that the Recycler stashed three times during the last 12 hours. Accelerator status: store 5411 was at 64.58E30, the Pbar stack was at 14.9E10 and the stash was at 168.8E10.

At 8:20 AM, LRF3 tripped off for the first of three times in about 90 minutes. It reset.

Operation reported at 9:56 AM that the antiproton production was down. With the help of a Pbar expert operators moved the target to a new location and production improved.

At 10 AM, LRF3 tripped off on a driver fault. Operators tried to reset the station and heard a loud bang. Operations called in experts. With LRF3 off, all downstream beam was also off.

Operations reported at 12:11 PM that earlier the Pelletron VME crate had hung up. Operators also found that the SF6 water skid had tripped off. Operators contacted experts and per instructions, restarted the water skid, restored a file for the VME, and then turned on the Pelletron.

At 5:47 PM, a Pbar Debuncher vacuum pump (IP801S) tripped off. Operations reset the pump, but it tripped off 20 minutes later. Operators left the pump off.

Linac experts reported at 6:44 PM that they were still working on the LRF3 amplifier.

At 7:40 PM, a Booster vacuum pump (IPRF23) tripped off and caused BRF13 & 14 to trip off. Everything reset okay.

Operations reported at 8:06 PM that there were no stashes during the last 12 hours due to trouble with LRF3. Accelerator status: store 5411 was at 36.53E30, the Pbar stack was at 42.0E10 and the stash was at 168.8E10.

At 10:06 PM, BRF14 tripped off on a screen power supply fault and would not reset. Operations bypassed the station.

Monday May 7

The midnight shift began with Operations monitoring store 5411 and the stash, with Pbar not stacking, and with NuMI and MiniBooNE not taking beam due to LRF3 problems

At 12:37 AM, Operations reported that LRF3 was up and running. A Linac expert reported that there were multiple failures that caused other failures in the station.

At 1:42 PM, LRF3 tripped off. It reset.

Operations terminated store 5411 at 3:17 AM.

Operations put the TeV into shot setup at 3:32 AM.

At 4:41 AM, LRF3 tripped off. It reset.

At 4:55 AM, KRF3 tripped off.

Operations established store 5412 at 5:43 AM with an initial luminosity of 150.31E30.

Operations established beam to MTest at 6:11 AM.

At 6:43 AM, LRF3 tripped off. It reset.

At 7:27 AM, BRF7 tripped off. It reset.

Operations reported at 7:48 AM that the Recycler stashed twice during the last 12 hours.

The Plans for Monday

The plans for today are to stack, stash, and store. The Run Coordinator reported that he would have to schedule interruptions to this plan in order to conduct necessary safety system test throughout this week. This schedule hasn’t been set yet.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 05/07/2007    email Fermilab